# Exploring Iterative Controllable Summarization with Large Language Models

### v2

@@ Abstract
Large language models (LLMs) have demonstrated remarkable performance in abstractive summarization tasks. However, their ability to precisely control summary attributes (e.g., length or topic) remains underexplored, limiting their adaptability to specific user preferences. In this paper, we systematically explore the controllability of LLMs. To this end, we revisit summary attribute measurements and introduce iterative evaluation metrics, failure rate and average iteration count to precisely evaluate controllability of LLMs, rather than merely assessing errors. Our findings show that LLMs struggle more with numerical attributes than with linguistic attributes. To address this challenge, we propose a guide-to-explain framework (GTE) for controllable summarization. Our GTE framework enables the model to identify misaligned attributes in the initial draft and guides it in self-explaining errors in the previous output. By allowing the model to reflect on its misalignment, GTE generates well-adjusted summaries that satisfy the desired attributes with robust effectiveness, requiring surprisingly fewer iterations than other iterative approaches.
@@
Sangwon Ryu1, Heejin Do1, Daehee Kim1, Hwanjo Yu12, Dongwoo Kim12, Yunsu Kim3, Gary Geunbae Lee12, Jungseul Ok12 1Graduate School of Artificial Intelligence, POSTECH, South Korea 2Department of Computer Science and Engineering, POSTECH, South Korea 3aiXplain Inc., Los Gatos, CA, USA {ryusangwon, heejindo, andrea0119, hwanjoyu, dongwoo.kim, gblee, jungseul}@postech.ac.kr, yunsu.kim@aixplain.com
@@ 1 Introduction
Large language models (LLMs) have demonstrated superior performance in text summarization, outperforming encoder-decoder models by generating more contextually appropriate and natural summaries Goyal et al. (2023); Zhang et al. (2024); Pu et al. (2023); Ryu et al. (2024b). However, given individuals’ diverse preferences for summary styles, it is essential to generate summaries tailored to individual needs Zhang et al. (2023b). For instance, some users may prefer specific topic-focused summaries or retain exact phrases. While LLMs excel in generating fluent summaries, their ability to precisely control attributes remains underexplored Liu et al. (2024), limiting their adaptability to diverse user preferences.
Controllable summarization has recently garnered attention Zhong et al. (2021); Ahuja et al. (2022); Maddela et al. (2022); Mehra et al. (2023); Xu et al. (2023); Zhang et al. (2023b), with prior studies primarily utilizing encoder-decoder architectures Mao et al. (2022); Zhang et al. (2022); Vig et al. (2022); He et al. (2022); Pagnoni et al. (2023); Wang et al. (2023); Urlana et al. (2024) that incorporate attribute-specific training signals to constrain summary generation. However, these systems lack scalability due to their dependence on attribute-specific training. In contrast, LLMs offer flexibility yet rely on in-context learning and prompting without explicit training-based constraints, hindering fine-grained attribute control Yuan et al. (2024); Tang et al. (2023), necessitating further investigation.
@@
To systematically explore the controllability of LLMs, we first revisit the measurements for four key attributes: extractiveness, length, topic, and speaker, and refine them for more precise measurement. Specifically, instead of previous attribute measurement strategies that rely solely on word presence for topic- or speaker-focused summaries, we adopt embedding similarity to incorporate semantic information into the measurements. Building on this refined attribute measurement, we move beyond the previous controllability evaluation, which primarily assesses the disparity between the reference and the generated summary He et al. (2022); Zhang et al. (2023b). To fully explore the LLMs’ controllability, we shift focus from merely measuring attribute errors to evaluating whether LLMs can accurately control specified attributes via iterative refinement. Thus, we introduce two evaluation metrics: the failure rate—the proportion of control failures within the maximum iterations—and the average iteration count required for successful control. Our findings reveal that while LLMs excel at controlling linguistic attributes such as topic and speaker, they severely struggle with numerical attributes such as extractiveness and length. We assume that unlike linguistic attributes, which rely on semantic coherence, numerical ones demand strict quantitative constraints, making fine-grained control challenging.
@@
To address this challenge, we propose a guide-to-explain (GTE) framework, which enables precise attribute control solely through LLMs without relying on additional attribute-specific training. We first design a step-by-step attribute identification phase to instruct the model on calculating misaligned attributes of its previously generated summary, then guide it to explain the rationales behind its errors. Through self-reflection, the model corrects its previous mistakes and generates a well-aligned summary in the regeneration phase. By integrating a self-refinement strategy—proven effective in complex reasoning tasks Weng et al. (2023); Madaan et al. (2023); Dhuliawala et al. (2024); Gou et al. (2024)—into controllable summarization, we enhance LLMs’ attribute controllability while ensuring summary quality.
@@
We evaluate GTE on mixed-attribute summarization datasets, MACSumDoc and MACSumDial Zhang et al. (2023b). GTE successfully controls each attribute with minimal iterations, significantly outperforming other iteration methods, while demonstrating robustness by consistently adjusting attributes across data samples. Further, we also demonstrate the high quality of the controlled summaries across multiple generic summarization evaluation metrics, including UniEval Zhong et al. (2022) and QuestEval Scialom et al. (2021). Finally, we analyze whether LLMs can control multiple attributes simultaneously, revealing their struggles in jointly managing correlated numerical attributes. Our contributions are as follows:
  • We systematically explore LLM’s controllability in text summarization.
@@
  • We refine the measurement of summarization attributes and introduce iterative evaluation metrics for evaluating LLMs’ controllability.
@@
  • We propose a guide-to-explain (GTE), which guides the model to explain its misalignments and effectively adjusts misaligned attributes within just a few iterations.
Attribute Metrics Paper
Extractiveness ROUGE, word overlap Goyal et al. (2022); Zhang et al. (2023b)
Length Absolute length, length ratio Goyal et al. (2022); He et al. (2022); Maddela et al. (2022); Zhang et al. (2023b)
Topic, Query ROUGE, LDA, topic word count, classifier Zhong et al. (2021); He et al. (2022); Zhang et al. (2023b); Xu et al. (2023)
Speaker, Entity ROUGE, speaker utterance word overlap Maddela et al. (2022); Zhang et al. (2023b)
Table 1: Previous methods for measuring attributes.
@@ 2 Related work
Controllable summarization has recently gained attention due to its practical applications Zhong et al. (2021); Ahuja et al. (2022); Maddela et al. (2022); Mehra et al. (2023); Xu et al. (2023); Zhang et al. (2023b); Ribeiro et al. (2023). Previous research employed encoder-decoder models to control attributes Fan et al. (2018); Liu and Chen (2021); Dou et al. (2021); He et al. (2022); Mao et al. (2022); Zhang et al. (2022); Goyal et al. (2022); Vig et al. (2022); Bahrainian et al. (2022); Liu et al. (2022); Pagnoni et al. (2023); Wang et al. (2023); Urlana et al. (2024). For example, CTRLSum He et al. (2022) trains models by prepending a keyword sequence to the source document. Similarly, MACSum Zhang et al. (2023b) adopts prompt learning by prepending each attribute’s value to the input source using a combination of hard prompts and soft prefixes. HYDRASUM Goyal et al. (2022) leverages a single encoder and multiple decoder framework, using a mixture-of-experts where the decoders share probabilities to effectively control the attributes.
Most controllable summarization research relied on encoder-decoder frameworks. In addition, they required attribute-specific training or custom datasets to control each attribute, limiting the flexibility of attribute manipulation. Therefore, we propose a generalizable approach to enable flexible attribute control without the need for tailored training for each attribute, leveraging underexplored LLMs for controllable summarization Tang et al. (2023); Yuan et al. (2024); Liu et al. (2024).
@@
Recently, the self-correction approach has been used to refine their initial attempts to solve complex problems Weng et al. (2023); Shinn et al. (2023); Madaan et al. (2023); Dhuliawala et al. (2024); Gou et al. (2024), mirroring humans. In summarization tasks, self-correction has been employed to enhance the overall quality of summaries Zhang et al. (2023a); Sun et al. (2024). Zhang et al. (2023a) utilizes iterative feedback from an evaluator to instruct ChatGPT to produce higher-quality summaries. Unlike prior works, we focus on generating summaries tailored to user preferences, which involve a multitude of factors to consider.
3 Problem formulation for controllable summarization
We conduct a systematic exploration of LLMs’ controllability evaluation. While generic evaluation metrics (e.g. consistency, fluency) favor higher scores, controllability requires aligning summaries with user-specified attribute values. Previous research mainly assessed deviations from target values rather than ensuring precise control Zhang et al. (2023b). To address this, we introduce iterative evaluation metrics, assessing whether LLMs can precisely adjust attributes, even through multiple iterations. Before evaluating a controllability, we first refine attribute measurement.
3.1 Controllable attribute measurement
We revisit attribute measurement to quantify key attributes for controllable summarization: extractiveness, length, topic, and speaker. Table 1 summarizes how previous controllable summarization studies have measured each attribute. However, the measurements for certain attributes have not yet been clearly defined. Thus, we outline our newly defined approach for attribute measurements below.
@@ 3.1 Controllable attribute measurement
quantifies the degree of lexical overlap between a summary and its source document. Highly extractive summary is required when users need to retain the original context, such as in legal documents, whereas paraphrasing is preferred in general cases. To measure the extractiveness, MACSum Zhang et al. (2023b) employs the average precision scores of ROUGE-2 and ROUGE-3. HYDRASUM Goyal et al. (2022) calculates the proportion of words in the summary that appear in the source document as well as the average length of copied contiguous spans in the summary. Following the definition of extractiveness, we measure this attribute as the proportion of words in the summary directly reused from the source text.
Attribute Label MACSumDoc MACSumDial
Distributions Relabel # of summaries Distributions Relabel # of summaries
Extractiveness normal 35.7 - 100.0% (85.2%) 85.0% 3731 53.2 - 100.0% (86.4%) 85.0% 1661
high 55.0 - 100.0% (90.0%) 90.0% 287 63.0 - 100.0% (88.9%) 90.0% 340
fully 84.6 - 100.0% (99.7%) 100.0% 260 75.9 - 100.0% (98.4%) 100.0% 337
Length short 0.7 - 15.0% (4.8%) 7.5% 1059 0.2 - 20.8% (2.0%) 7.5% 300
normal 0.5 - 48.6% (6.9%) 20.0% 2194 0.3 - 41.9% (3.7%) 20.0% 1693
long 1.5 - 39.8% (13.9%) 32.5% 1025 0.7 - 32.4% (6.0%) 32.5% 345
Topic - 74.8 - 88.8 74.0 2013 73.6 - 87.0 74.0 2317
Speaker - - - - 75.6 - 92.0 75.0 1796
Table 2: Data distributions of MACSumDoc and MACSumDial.
@@
refers to the number of words or sentences in the summary or the ratio of the summary’s length to the original text. By controlling the length, the amount of information in the summary can be adjusted according to user preferences. Prompts suggested by earlier works specify a fixed number of sentences, e.g., "3 sentences," but this approach fails to account for variations in sentence length and does not accurately reflect the summary’s actual length Goyal et al. (2023); Liu et al. (2024); Yuan et al. (2024). As the summary length may vary depending on the source document, we use the length ratio instead of absolute length.
@@
refers to generating a summary centered around one or multiple themes. Query-focused summarization (QFS), which generates summaries based on a specific query, and entity-based summarization, which focuses on a particular individual, are variations of topic-focused summarization. Zhang et al. (2023b) measured topic word frequency in summaries. Similarly, most QFS methods have relied solely on ROUGE scores, evaluating generated summaries by comparing them to human-annotated references Zhong et al. (2021). However, even if topic words do not explicitly appear, the summary can still reflect the core context of the topic, especially for LLM-generated summaries, which tend to paraphrase the content. Therefore, rather than simply counting the frequency of word occurrences, we evaluate the semantic similarity between the summary and each topic-related word. We compute the embedding similarity $\mathcal{B}$ between the topic word and each word in the summary $s$: $\frac{1}{n}\sum_{i\in s}\mathcal{B}(topic,word_{i})$, where $n$ is the number of words in the summary. If multiple topics $k$ are present, we use the average embedding similarity across all topics: $\frac{1}{k}\sum_{j\in k}\frac{1}{n}\sum_{i\in s}\mathcal{B}(topic_{k},word_{i})$.
@@
refers to generating a summary that focuses on the utterances of a specific speaker from a long document or dialogue. Zhang et al. (2023b) calculate the frequency of the speaker’s spoken words appearing in the summary. Similar to topics, measuring the proportion of words from a specific speaker’s dialogue included in the summary does not fully capture semantic elements. Therefore, we extract speaker utterances and construct speaker set $\mathcal{U}$ and leverage BERTScore F1 Zhang et al. (2020) to compute the embedding similarity between the summary $s$ and $\mathcal{U}$: BERTScore$(s,\mathcal{U})$.
3.2 Iterative controllablility evaluation
@@
We set attribute-specific thresholds and iteratively regenerate summaries until they are met. Each attribute is measured using the criteria outlined in Section 3.1 to determine its threshold. For extractiveness and length, we consider control successful if the attribute values fall within $\pm$5 of the target value. For topic and speaker, we use the minimum embedding similarity values of the reference summaries in the training dataset (Table 2) as thresholds to determine whether a summary is topic-focused or speaker-focused. The threshold can be adjusted based on the strictness of the evaluation criteria.
Label reinterpretation
We use the two publicly available MACSum datasets Zhang et al. (2023b) for controllable summarization. However, existing labels are ambiguous since the criteria are not numerically defined (e.g., how short must a summary be to qualify as short?). We believe that these ambiguous criteria may confuse LLMs, so we assign clear numerical values to each label. To provide detailed criteria, we reinterpret the labels based on the attribute distributions in each training set. For extractiveness, we set labels as {normal: 85%, high: 90%, fully: 100%}, following the mean value. Unlike Zhang et al. (2023b), we define the summary length as a ratio of the original text rather than a fixed value. The existing labels do not distinctly differentiate these ratios, as the average length ratios between labels show minimal differences. Specifically, short is 2.0, while normal is 3.7, indicating a relatively small gap. To allow for greater variability, we expand the range. Importantly, our method generates summaries based on the specified numerical values, regardless of predefined labels. The broader range allows for a more adaptive and effective evaluation of LLM controllability. we set the lengths to {short: 7.5%, normal: 15%, long: 32.5%}, providing clearer distinctions with evaluating a broader range of controllability.
[…]
4 Controllability of LLMs in text summarization
@@
Table 3: The left number represents the averaged control failure rate, and the right side denotes the average iteration count for successful control.
As research on leveraging LLMs for controllable summarization remain limited, we evaluate the controllability of various LLMs using the iterative evaluation method from Section 3.2. We first provide an initial control prompt and generate a summary. If the generated summary fails to meet the specified attribute threshold, the result is fed back into the LLM’s input, prompting it to regenerate until the attribute is correctly controlled. As shown in Table 3, smaller-scale LLMs, such as Phi-3-medium Abdin et al. (2024) and Llama3-8B Dubey et al. (2024), partially control topic, but fail to control extractiveness and length. Similarly, large-scale LLMs such as Llama3-70B, GPT-3.5 Brown (2020), and GPT-4o Achiam et al. (2023) effectively control topic, demonstrating low failure rates. However, they also struggle to control extractiveness and length, showing failure rates of around 50%. Notably, when initial attempts fail, GPT-3.5 and GPT-4o could not adjust despite multiple iterations, ultimately reaching the maximum iteration limit, resulting in an iteration count of zero. As a result, we find that generating summaries while controlling attributes remains challenging for LLMs, even with iterative attempts, especially for numerical attributes such as extractiveness and length.
Model Extractiveness ($\downarrow$ / $\downarrow$) Length ($\downarrow$ / $\downarrow$) Topic($\downarrow$) Speaker($\downarrow$)
normal high full avg short normal long avg

### v3

@@ Abstract
Large language models (LLMs) have demonstrated remarkable performance in abstractive summarization tasks. However, their ability to precisely control summary attributes (e.g., length or topic) remains underexplored, limiting their adaptability to specific user preferences. In this paper, we systematically explore the controllability of LLMs. To this end, we revisit summary attribute measurements and introduce iterative evaluation metrics, failure rate and average iteration count, to more precisely evaluate controllability beyond assessment of errors. Our findings show that LLMs struggle more with numerical attributes than with linguistic attributes. To address this challenge, we propose a guide-to-explain framework (GTE) for controllable summarization. GTE enables the model to identify misaligned attributes in the initial draft and guides it to self-explain errors in the previous output. By encouraging reflection on attribute misalignment, GTE generates well-adjusted summaries that satisfy the desired attributes with robust effectiveness while requiring surprisingly fewer iterations than other iterative approaches.
@@
Sangwon Ryu1, Heejin Do3, Daehui Kim4, Hwanjo Yu1,2, Dongwoo Kim1,2, Yunsu Kim5, Gary Geunbae Lee1,2, Jungseul Ok1,2 1GSAI, POSTECH  2CSE, POSTECH 3ETH AI Center  4AI future lab, KT  5LILT {ryusangwon, hwanjoyu, dongwoo.kim, gblee, jungseul}@postech.ac.kr heejin.do@ai.ethz.ch daehui.kim@kt.com yunsu.kim@lilt.com
@@ 1 Introduction
Large language models (LLMs) have demonstrated superior performance in text summarization, outperforming encoder-decoder models by generating more contextually appropriate and natural summaries Goyal and others (2023); Zhang et al. (2024); Pu et al. (2023); Ryu et al. (2024a). However, given the diversity of individual preferences for summary styles, it is essential to generate summaries tailored to specific user needs Zhang et al. (2023b). For example, some users may prefer topic-focused summaries or wish to retain exact phrases. Although LLMs excel at generating fluent summaries, their ability to precisely control attributes remains underexplored Liu et al. (2024), limiting their adaptability to diverse user preferences. Typical requests can be ambiguous, such as “summarize in 3 sentences” or “generate a highly extractive summary”. Sentence lengths can vary significantly, and vague terms such as “highly” hinder reliable instruction-following and complicate evaluating whether the instructions are properly satisfied (Figure 1).
Figure 1: Ambiguous instructions hinder LLMs’ ability to follow control signals and complicate the evaluation process (e.g., how should “highly” be judged in a generated summary?).
@@
Therefore, we systematically explore the controllability of LLMs. We begin by revisiting the measurements for four key attributes: extractiveness, length, topic, and speaker, and refine them for more precise measurement. For length, instead of dividing summaries into fixed-length buckets and computing differences across them, which overlooks variations within each bucket, we directly measure the gap between the generated and target length. Moreover, rather than relying solely on word presence as in previous strategies for measuring topic- or speaker-focused summaries, we adopt embedding-based similarity to incorporate semantic information into the measurements. With more precise attribute measurements in place, we next investigate how reliably LLMs can control these attributes. To fully explore LLM controllability, we evaluate whether LLMs can accurately control specified attributes through iterative refinement. Even if initial attempts fail, we test whether they can eventually succeed without external guidance. To this end, we introduce two evaluation metrics: the failure rate—the proportion of control failures within the maximum iterations—and the average iteration count until successful control. In Section 4, we show that while LLMs excel at controlling linguistic attributes such as topic and speaker, they struggle significantly with numerical attributes such as extractiveness and length. We assume that, unlike linguistic attributes, which rely on semantic coherence, numerical ones require adherence to strict structural constraints, making fine-grained control challenging.
@@
To address this challenge, we introduce a guide-to-explain (GTE) framework, a self-correction-based approach that enables precise attribute control through LLMs without relying on additional attribute-specific training. Although self-correction approaches have been employed to evaluate and revise their own outputs for complex tasks, we observe that LLMs often fail to reliably measure their output summary attributes. Therefore, we provides the model with guidance and encourages it to reflect on why its previous attempt was incorrect before regenerating the summary. We first design a step-by-step attribute identification phase that instructs the model to identify misaligned attributes in its previously generated summary and then guides it to explain the rationale behind its errors. Through self-reflection, the model corrects its prior mistakes and generates a well-aligned summary in the regeneration phase. By integrating a self-refinement strategy into controllable summarization, we improve the controllability of LLMs while enhancing summary quality.
@@
We evaluate GTE on mixed-attribute summarization datasets, MACSumDoc and MACSumDial Zhang et al. (2023b). Whereas standard iterative methods often prove ineffective for controllable summarization, our modified approach achieves reliable attribute control with minimal iterations and consistently adjusts attributes across data samples. Furthermore, we demonstrate the high quality of the controlled summaries across multiple generic summarization evaluation metrics, including UniEval Zhong et al. (2022) and QuestEval Scialom et al. (2021). Finally, we analyze whether LLMs can control multiple attributes simultaneously, revealing their difficulty in jointly managing correlated numerical attributes. Our contributions are as follows:
  • We systematically explore the controllability of LLMs in text summarization.
@@
  • We refine the measurement of summarization attributes and introduce iterative evaluation metrics to evaluate LLM controllability.
@@
  • We propose a guide-to-explain (GTE) framework, which guides the model to explain its misalignments and effectively adjust attributes within just a few iterations.
@@ 2 Related work
Controllable summarization has recently gained attention due to its practical applications Zhong et al. (2021); Ahuja et al. (2022); Maddela et al. (2022); Mehra et al. (2023); Xu et al. (2023); Zhang et al. (2023b); Ribeiro et al. (2023); Retkowski and Waibel (2025). Previous research has employed encoder-decoder models to control attributes Fan et al. (2018); Liu and Chen (2021); Dou et al. (2021); He et al. (2022); Mao et al. (2022); Zhang et al. (2022); Goyal et al. (2022); Vig et al. (2022); Bahrainian et al. (2022); Liu et al. (2022); Pagnoni et al. (2023); Wang et al. (2023); Urlana et al. (2024). For example, CTRLSum He et al. (2022) trains models by prepending a keyword sequence to the source document. Similarly, MACSum Zhang et al. (2023b) adopts prompt learning by prepending each attribute’s value to the input using a combination of hard prompts and soft prefixes. HYDRASUM Goyal et al. (2022) leverages a single encoder, multiple decoder framework with a mixture-of-experts approach, where the decoders share probabilities to effectively control the attributes.
Most controllable summarization research has relied on encoder-decoder frameworks. In addition, these methods often require attribute-specific training or custom datasets to control each attribute, limiting the flexibility of attribute manipulation. Therefore, we propose a generalizable approach that enables flexible attribute control without the need for tailored training, leveraging LLMs for controllable summarization Tang et al. (2023); Yuan et al. (2024); Liu et al. (2024).
@@
Recently, self-correction approaches have been used to refine initial attempts at solving complex problems Weng et al. (2023); Shinn et al. (2023); Madaan et al. (2023); Dhuliawala et al. (2024); Gou et al. (2024), mirroring human behavior. In summarization tasks, self-correction has been employed to enhance the overall quality of summaries Zhang et al. (2023a); Sun et al. (2024). Zhang et al. (2023a) utilizes iterative feedback from an evaluator to instruct ChatGPT to produce higher-quality summaries. Unlike prior work, we focus on generating summaries tailored to user preferences, which involve multiple factors to consider.
3 Attribute Measurement and Evaluation Framework for LLM Controllability
We first analyze how each summarization attribute has traditionally been measured and redefine those that were not clearly defined. In particular, we refine linguistic attributes—often measured by word count, using embedding-based similarity. These refined measurements allow us to more accurately capture the attributes of generated summaries. Building on this, we propose iterative evaluation metrics to assess the controllability of LLMs—that is, their ability to precisely adjust attributes through multiple rounds of control.
3.1 Revisiting attribute measurements for controllable summarization
We revisit attribute measurement to quantify key attributes for controllable summarization: extractiveness, length, topic, and speaker. We observe that the measurements for certain attributes have not yet been clearly defined, leading to unclear criteria for evaluating controllability. Therefore, we outline our newly defined approach for attribute measurements below.
@@ 3.1 Revisiting attribute measurements for controllable summarization
quantifies the degree of lexical overlap between a summary and its source document. a highly extractive summary is preferred when users need to retain the original context, such as in legal documents, whereas paraphrasing is often favored in general cases. Following the definition of extractiveness, we measure the attribute as the proportion of words in the summary directly reused from the source text.
@@
refers to the number of words or sentences in the summary or the ratio of the summary’s length to that of the original text. By controlling the length, the amount of information in the summary can be adjusted according to user preferences. Prompts used in earlier work often specify a fixed number of sentences (e.g., "3 sentences"), but this approach fails to account for variations in sentence length and does not accurately reflect the summary’s actual length Goyal and others (2023); Liu et al. (2024); Yuan et al. (2024). Since summary length may vary with document complexity Ryu et al. (2024b) and the original dataset Zhang et al. (2023b) annotates length based on ratios, we use the length ratio rather than absolute length in our experiments. Further, prior studies He et al. (2022); Zhang et al. (2023b) evaluated length controllability by grouping outputs into discrete buckets and measuring differences across buckets. However, this approach overlooks variations within each bucket. Thus, we directly compare the generated summary length with the requested target length.
@@
refers to generating a summary centered around one or more themes. Query-focused summarization (QFS), which generates summaries based on a specific query, and entity-based summarization, which focuses on a particular individual, are variations of topic-focused summarization. Most prior work has measured topic word frequency in summaries Fan et al. (2018); He et al. (2022); Zhang et al. (2023b). Similarly, most QFS methods have relied solely on ROUGE scores, evaluating generated summaries by comparing them to human-annotated references Zhong et al. (2021). However, even when topic words do not explicitly appear, a summary can still reflect the core context of the topic—especially in LLM-generated summaries, which tend to paraphrase content. Therefore, rather than simply counting word occurrences, we evaluate the semantic similarity between the summary and each topic-related word. We compute the embedding similarity $\mathcal{B}$ between the topic word and each word in the summary $s$ as follows: $\frac{1}{n}\sum_{i\in s}\mathcal{B}(topic,word_{i})$, where $n$ is the number of words in the summary. If multiple topics $k$ are present, we use the average embedding similarity across all topics: $\frac{1}{k}\sum_{j\in k}\frac{1}{n}\sum_{i\in s}\mathcal{B}(topic_{j},word_{i})$.
@@
3.2 Iterative controllability evaluation
Building on these refined measurements of summary attributes, we introduce iterative evaluation metrics to assess whether LLMs can iteratively refine and adjust attributes over multiple revisions. Specifically, we introduce two metrics: (1) the failure rate, proportion of cases in which the model reaches the predefined maximum number of iterations without achieving the desired modifications, and (2) the average iteration count required for successful attribute control. We set the maximum number of iterations to 20 due to cost constraints.
@@ 3.2 Iterative controllability evaluation
We set attribute-specific thresholds and iteratively regenerate summaries until those thresholds are met. Each attribute is measured using the criteria outlined in Section 3.1 to determine its respective threshold. For extractiveness and length, we consider control successful if the attribute values fall within $\pm$5 of the target value. For topic and speaker, we use the minimum embedding similarity values of the reference summaries in the training dataset as thresholds to determine whether a summary is topic-focused or speaker-focused. Theses thresholds can be adjusted based on the strictness of the evaluation criteria. The distribution of the datasets used in our experiments is provided in Appendix A.
@@
Table 1: We evaluate the controllability of LLMs by iteratively testing their ability to accurately adjust specified attributes. The left number represents the averaged control failure rate, and the right side denotes the average iteration count for successful control.
Label reinterpretation
We use the two publicly available MACSum datasets Zhang et al. (2023b) for controllable summarization. However, existing labels are ambiguous, as the criteria are not numerically defined (e.g., how short must a summary be to qualify as short?). We believe that such ambiguity may confuse LLMs, so we assign clear numerical values to each label. To provide detailed criteria, we reinterpret the labels based on the attribute distributions in each training set (see Appendix A). For extractiveness, we define the labels as normal: 85%, high: 90%, and fully: 100%, based on the average values. For the length attribute, we follow the annotation criteria of the MACSum dataset—short: 5–10%, normal: 15–25%, and long: 30–35%—and set our target values to short: 7.5%, normal: 15%, and long: 32.5%. Importantly, our method generates summaries based on the specified numerical values, regardless of predefined labels.
4 Analysis on Controllability of LLMs
4.1 Iterative Evaluation on LLMs
Figure 2: LLMs show notable errors in word count estimation: for an article with 484 words and a summary with 157 words, the model predicts 668 and 159 words, respectively—revealing limitations in self-critique within controllable summarization.
[…]
As research on leveraging LLMs for controllable summarization remains limited, we evaluate the controllability of various LLMs using the iterative evaluation method described in Section 3.2. We first provide an initial control prompt and generate a summary. If the generated summary fails to meet the specified attribute threshold, the result is fed back into the LLM’s input, prompting it to regenerate until the attribute is correctly controlled. As shown in Table 1, smaller-scale LLMs such as Phi-3-medium Abdin et al. (2024) and Llama3-8B Dubey et al. (2024), partially control topic, but fail to control extractiveness and length. Similarly, large-scale LLMs such as Llama3-70B, GPT-3.5 Brown (2020), and GPT-4o Achiam et al. (2023) effectively control topic, demonstrating low failure rates. However, they struggle with extractiveness and length, with failure rates of around 50%. Notably, when initial attempts fail, even GPT-4o is unable to adjust after multiple iterations, ultimately reaching the maximum iteration limit, resulting in an iteration count of zero. These findings suggest that generating summaries while controlling attributes remains challenging for LLMs, even with iterative attempts, especially for numerical attributes such as extractiveness and length.
4.2 Self-correction for controllable summarization
We evaluate whether LLMs can adjust summary attributes through self-correction, which has previously improved performance in generic summarization Zhang et al. (2023a); Sun et al. (2024). However, unlike in generic summarization tasks, LLMs struggle to measure attributes, as shown in Figure 2. Specifically, they fail to accurately count words in either the source or the summary, making it infeasible for them to revise summaries to match target attribute values on their own.
5 Guide-to-Explain (GTE)
Therefore, we introduce a guide-to-explain (GTE) framework to control attributes via LLMs. As shown in Figure 3, the GTE consists of two key phases: step-by-step attribute identification and self-explanation guidance. Since LLMs struggle to reliably measure summary attributes on their own (Figure 2), we explicitly provide the attribute values and teach the model step by step how each attribute should be identified. We then guide the LLM to reflect by explaining the rationale behind its mistakes, enabling it to make appropriate adjustments in subsequent iterations.
Model Extractiveness ($\downarrow$ / $\downarrow$) Length ($\downarrow$ / $\downarrow$) Topic($\downarrow$ / $\downarrow$)
normal high fully avg short normal long avg